Solve for the electric flux of a sphere if the radius r of the sphere is 10.57 nm and its charge q is 6.11 nC


1
Expert's answer
2022-05-22T14:39:07-0400

The Gauss' law says

"\phi=\oint EdA=\frac{q}{\epsilon_0}"

"\phi=\frac{6.11*10^{-9}}{8.85*10^{-12}}=690\:\rm N\cdot m^2/C"